Let X1N(4.2,1) and X2~ N(12,70 be two Gaussian random variables.

(1) Sketch the PDFs of X_{1} , X_{2} on the same chart.

(2) Assuming X_{1} , X_{2} are independent, compute mathbb{P}[X_{1}>X_{2}] .
Answer #1

Histogram of x 4 Histogram of y -10 10 20 30 40

code in R

par(mfrow = c(2,1))


x <- rnorm(1000, mean=4.2, sd=1)
hist(x, probability=TRUE)
xx <- seq(min(x), max(x), length=100)
lines(xx, dnorm(xx, mean=4.2, sd=1))

y <- rnorm(1000, mean=12, sd= sqrt(70))
hist(y, probability=TRUE)
yy <- seq(min(y), max(y), length=100)
lines(yy, dnorm(yy, mean=12, sd= sqrt(70)))

2)

x1 = N( 4.2 , 1)

X2 = N(12,70)

X1 - X2 = N(4.2 - 12 , 1 + 70)

= N(-7.8 , 71)

P(X1 > X2)

= P(X1 - X2 > 0)

= P(Z > (0 -(-7.8))/sqrt(71))

=P(Z > 0.92568 )

=

0.1773062